The import duty on a certain electronic product increased by 20%. the demand for it thereby reduced by 10% is there a percentage increase or decrease in revenue and by how much?
1
Expert's answer
2019-11-01T15:08:42-0400

"Q_0 = K_0*P" is the revenue before changes.


After increasing the import duty by 20% the product price will be "P+0.2P=1.2P" and after a reduction the demand will be "K_0-0.1K_0=0.9K_0" .


After changing the revenue is "Q_n = 1.2P*0.9K_0 = 1.08*K_0*P=1.08Q_o" .

The difference between the revenues after changes and before changes is "Q_n-Q_0=1.08Q_0-Q_0=0.08Q_0" .

It means that increase in revenue is 8%.


Answer: increase in revenue is 8%.